# Subtleties in the pseudomodes formalism

### v1

@@ Abstract
The pseudomode method for open quantum systems, also known as the mesoscopic leads approach, consists in replacing a structured environment by a set of auxiliary “pseudomodes” subject to local damping that approximate the environment’s spectral density. Determining what parameters and geometry to use for the auxiliary modes, however, is non-trivial and involves many subtleties. In this paper we revisit this problem of pseudomode design and investigate some of these subtleties. In particular, we examine the scenario in which pseudomodes couple to each other, resulting in an effective spectral density that is no longer a sum of Lorentzians. We show that non-diagonalizability of the pseudomodes’ effective single-particle non-Hermitian Hamiltonian can lead to terms in the effective spectral density which cannot be obtained by diagonalizable non-Hermitian Hamiltonians. We also present a method for constructing the pseudomode parameters to exactly match a fit to a spectral density, and in doing so illuminate the enormous freedom in this process. The case of many uncoupled pseudomodes is explored, and we show how, contrary to conventional assumption, the effective spectral density does not necessarily converge in the limit of an infinite number of pseudomodes; we attribute this to the non-completeness of Lorentzians as basis functions. Finally, we discuss how the notion of effective spectral densities can also emerge in the context of scattering theory for non-interacting systems.
@@ I Introduction
The method has been used to treat problems in areas including fermionic transport Chen et al. (2014); Rams and Zwolak (2020); Wójtowicz et al. (2021); Elenewski et al. (2021); Brenes et al. (2023); Cirio et al. (2023), quantum thermodynamics Lacerda et al. (2023, 2024); Bettmann et al. (2025); Menczel et al. (2024); Albarelli et al. (2024), and quantum optics Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); Lednev et al. (2024); Liang et al. (2024). Recent works have combined this method with tensor networks Brenes et al. (2020); Cirio et al. (2023); Lacerda et al. (2024) and the Time Evolving Density operator with Orthogonal Polynomials Algorithm (TEDOPA) Nüßeler et al. (2022); Ferracin et al. (2024) to treat interacting systems. One advantage of this method is that by modeling the environments using physical modes to form an extended system, we have access to details about correlations between the system and its environment at different energies. This has been used to gain insight into energy-resolved currents for interacting fermionic systems Brenes et al. (2020). Likewise, using the framework of continuous measurement, one can monitor quantum jumps between the residual baths and pseudomodes and perform trajectory unraveling to probe stochastic thermodynamic quantities Bettmann et al. (2025). The auxiliary modes can therefore possess physical significance, rather than serving merely as a mathematical construct for modeling the environment. Alternatively, one can relax the constraint that the pseudomodes have physical parameters and include non-Hermitian terms in the pseudomode Hamiltonian or negative damping rates in the GKSL master equation in order to model an even wider variety of environments Lambert et al. (2019); Luo et al. (2023); Menczel et al. (2024).
@@
The success of the pseudomodes approach, however, relies on the careful design of their geometry and parameters so as to properly mimic the real environment. Implementing this entails various subtleties, which are often overlooked. The goal of this paper is to address said subtleties. In section II, we consider the effective spectral density that arises from a particular configuration of pseudomodes. Depending on the non-Hermitian Hamiltonian generated by the pseudomodes’ Hamiltonian and the residual baths, three possible cases emerge. If none of the pseudomodes couple to each other, then one obtains an effective spectral density that is a sum of Lorentzians, while if the pseudomodes couple to each other there are additional terms in the effective spectral density which we call “Anti-Lorentzian”. We take this one step further, and show that one can obtain terms which are neither Lorentzian nor Anti-Lorentzian by using a non-diagonalizable non-Hermitian Hamiltonian for the pseudomodes. In section III we discuss the reverse process of choosing pseudomode parameters to obtain a particular effective spectral density, which is required to actually implement the pseudomode method. This is generally done in two steps. First, the true spectral density is fit in Fourier space with a set of decaying complex exponentials. Then, the pseudomode parameters must be chosen to match that exponential fit, which is typically done by imposing some additional structure on the pseudomode parameters and then performing optimization Mascherpa et al. (2020). This second step is the greatest challenge to implementing the pseudomode method, in large part because there are many more free parameters for the pseudomodes than there are constraints given by the effective spectral density. We present a new method for choosing the pseudomode parameters, by which the parameters can be chosen to exactly match a decaying exponential fit. Our method also sheds light on the vast number of free parameters involved when choosing the pseudomode parameters. In section IV, we briefly explore the use of uncoupled pseudomodes uniformly distributed in energy and show that, contrary to conventional assumption, the effective spectral density obtained from this approach approximates but does not converge to the true bath spectral density. Instead, we present an alternate method that, while also not convergent in the infinite limit, significantly reduces the error between the true and effective spectral densities in the limit of many pseudomodes and is equally easy to implement. Finally, in section V we show how the effective spectral densities of section II also emerge in the context of scattering theory for non-interacting systems Blanter and Büttiker (2000); Büttiker (1992); Levitov et al. (1996); Esposito et al. (2009). We confirm that if the pseudomode parameters are chosen such that the effective spectral density exactly matches the true spectral density then the currents obtained from Landauer-Büttiker theory will also match exactly between the original system-bath description and pseudomode description. This also presents a new way to study scattering theory currents by breaking down a current between two real baths into the sum of currents between residual baths under a pseudomode mapping.
@@ I.1 Model System
$H_{\rm tot}=$ $H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k}\omega_{k}d_{\alpha k}^{\dagger}d_{\alpha k}$ (1)
$+\sum_{\alpha\in\{L,R\}}\sum_{i,k}g_{\alpha i,k}\big(c_{i}^{\dagger}d_{\alpha k}+d_{\alpha k}^{\dagger}c_{i}\big)\,$,
where $d_{\alpha k}$ are the fermionic modes of bath $\alpha\in\{L,R\}$. We set $\hbar=1$ here and throughout the paper. The system Hamiltonian can be completely general and may include arbitrary driving. The last term describes how bath $\alpha\in\{L,R\}$ can, in principle, couple to any mode $c_{i}$ of the system. Each bath is characterized by a spectral density matrix
$(J_{\alpha})_{ij}=2\pi\sum_{k}g_{\alpha i,k}g_{\alpha j,k}\delta(\omega-\omega_{k})\,$, (2)
@@
$\frac{dc_{i}(t)}{dt}=i[H_{S}(t),c_{i}(t)]-i\sum_{\alpha k}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ (3)
$-\sum_{\alpha kj}\intop_{0}^{t}dt^{\prime}g_{\alpha i,k}g_{\alpha j,k}e^{-i\omega_{\alpha k}(t-t^{\prime})}c_{j}(t^{\prime})\,$,
@@
$\frac{dc_{i}(t)}{dt}=i[H_{S}(t),c_{i}(t)]+\sum_{\alpha}\xi_{\alpha i}(t)-\sum_{\alpha j}\intop_{0}^{t}dt^{\prime}\chi_{\alpha,ij}(t-t^{\prime})c_{j}(t^{\prime})$ (4)
where $\xi_{\alpha i}=-i\sum_{k}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ is the noise operator, and the memory kernel is given by the Fourier transform of the spectral density:
@@
$C^{+}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)f_{\alpha}(\omega)$ (6)
@@
$\sum_{\alpha}\sum_{i}c_{i}^{\dagger}B_{\alpha i}+B_{\alpha i}^{\dagger}c\,$, (8)
@@
$C^{+}_{\alpha ij}(t)=\left\langle B_{i}^{\dagger}(t)B_{j}(0)\right\rangle\,$, (9)
$C^{-}_{\alpha ij}(t)=\left\langle B_{i}(t)B_{j}^{\dagger}(0)\right\rangle\,.$ (10)
@@ I.2 Pseudomode Mapping
$H_{\rm tot}=H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k,q}\Lambda_{\alpha,kq}a_{\alpha k}^{\dagger}a_{\alpha q}$
$\qquad+\sum_{\alpha\in\{L,R\}}\sum_{i,k}\big(\zeta_{\alpha i,k}a_{\alpha k}^{\dagger}c_{i}+\zeta_{\alpha i,k}^{*}c_{i}^{\dagger}a_{\alpha k}\big)$ (11)
$+\sum_{\alpha\in\{L,R\}}\sum_{k}\sum_{p}\Big\{\omega_{p}b_{\alpha k,p}^{\dagger}b_{\alpha k,p}+\tau_{\alpha k,p}\big(a_{\alpha k}^{\dagger}b_{\alpha k,p}+b_{\alpha k,p}a_{\alpha k}\big)\Big\}\,.$
The first line of (I.2) describes the possible interactions between pseudomodes of the same bath, described by the $n_{\alpha}\times n_{\alpha}$ matrix $\Lambda_{\alpha}$ for each bath. If $\Lambda$ is diagonal then none of the pseudomodes couple to each other, and we have the simpler configuration shown in Fig. 1(b). The second line describes how each pseudomode interacts with the central system through a set of $n_{\alpha}\times n_{S}$ matrices $\zeta_{\alpha}$. Finally, the last line describes a set of infinite dimensional residual reservoirs, one for each pseudomode $a_{\alpha k}$, composed of a set of modes $b_{\alpha k,p}$. The residual baths then have spectral densities given by
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p}\left|\tau_{\alpha k,p}\right|^{2}\delta(\omega-\omega_{p})\,$, (12)
which is a scalar, rather than a matrix like (2), since each residual bath $(\alpha,k)$ couples only to the pseudomode $a_{\alpha k}$. Generally, we want the couplings to the residual baths $\tau_{\alpha k,p}$ to be small compared to the extended system energy levels and constant at all energies, so that the residual baths have constant spectral densities
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha k}\,.$ (13)
In this case, the system feels an effective memory kernel given by Mascherpa et al. (2020):
@@
The pseudomode parameters are then chosen so that the effective spectral density (16) and effective memory kernel (14) are good approximations for the true spectral density (2) and memory kernel (5) of the original system. The derivation for these forms of $J^{\rm eff}_{\alpha,ij},\chi_{\alpha,ij}^{\rm eff}$ can be found in Appendix A. The reduced system dynamics under (1) will then be identical to those under (I.2) so long as the correlation functions from $J^{\rm eff}_{\alpha,ij}$ are identical to the true correlation functions (6); likewise, the dynamics under (1) will be well-approximated by the dynamics under (I.2) so long as the effective correlation functions well-approximate the true ones Trivedi et al. (2021); Huang et al. (2024). In practice, we achieve this by ensuring that $J^{\rm eff}_{\alpha,ij}(\omega)$ is a good fit to $J_{\alpha,ij}(\omega)$.
The utility of this method thus hinges on our ability to choose a set of $\zeta_{\alpha i}$, $\Lambda_{\alpha}$, and $\Gamma_{\alpha}$ so that the effective spectral density (16) closely matches the actual spectral density (2). In the next section, we describe how to obtain the form of $J^{\rm eff}_{\alpha,ij}$ from the chosen pseudomode configuration; in the following section, we describe the more complicated problem of choosing pseudomode parameters to obtain a desired $J^{\rm eff}_{\alpha,ij}$. As we shall show shortly, there are many possible choices of pseudomode parameters that give rise to the same effective spectral density, and not all of them may be equally desirable. Actually computing the extended-system dynamics generally requires further approximations, most commonly the use of a local GKSL master equation, which may be more or less reasonable to make based on the choice of pseudomode parameters. We do not discuss the merits of particular choices of parameters in this paper, instead focusing on the procedure for constructing the pseudomode parameters and emphasizing the freedoms therein.

### v2

@@ Abstract
The pseudomode method for open quantum systems, also known as the mesoscopic leads approach, consists in replacing a structured environment by a set of auxiliary “pseudomodes” subject to local damping that approximate the environment’s spectral density. Determining what parameters and geometry to use for the auxiliary modes, however, is non-trivial and involves many subtleties. In this paper we revisit this problem of pseudomode design and investigate some of these subtleties. In particular, we examine the scenario in which pseudomodes couple to each other, resulting in an effective spectral density that is no longer a sum of Lorentzians. We show that non-diagonalizability of the pseudomodes’ effective single-particle non-Hermitian Hamiltonian can lead to terms in the effective spectral density which cannot be obtained by diagonalizable non-Hermitian Hamiltonians. We also present a method for constructing the pseudomode parameters to exactly match a fit to a spectral density, and in doing so illuminate the enormous freedom in this process. The case of many uncoupled pseudomodes evenly distributed in energy is explored, and we show how, contrary to conventional assumption, the effective spectral density does not necessarily converge in the limit of an infinite number of pseudomodes distributed this way. Finally, we discuss how the notion of effective spectral densities can also emerge in the context of scattering theory for non-interacting systems.
@@ I Introduction
The method has been used to treat problems in areas including fermionic transport Chen et al. (2014); Schwarz et al. (2016); Elenewski et al. (2017); Rams and Zwolak (2020); Wójtowicz et al. (2021); Elenewski et al. (2021); Brenes et al. (2023); Cirio et al. (2023), quantum thermodynamics Lacerda et al. (2023, 2024); Bettmann et al. (2025); Menczel et al. (2024); Albarelli et al. (2024), quantum optics Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); Lednev et al. (2024); Liang et al. (2024), and strongly-correlated systems Arrigoni et al. (2013); Dorda et al. (2014, 2015). Recent works have combined this method with tensor networks Brenes et al. (2020); Cirio et al. (2023); Lacerda et al. (2024) and the Time Evolving Density operator with Orthogonal Polynomials Algorithm (TEDOPA) Nüßeler et al. (2022); Ferracin et al. (2024) to treat interacting systems. One advantage of this method is that by modeling the environments using physical modes to form an extended system, we have access to details about correlations between the system and its environment at different energies. This has been used to gain insight into energy-resolved currents for interacting fermionic systems Brenes et al. (2020). Likewise, using the framework of continuous measurement, one can monitor quantum jumps between the residual baths and pseudomodes and perform trajectory unraveling to probe stochastic thermodynamic quantities Bettmann et al. (2025). The auxiliary modes can therefore possess physical significance, rather than serving merely as a mathematical construct for modeling the environment. Alternatively, one can relax the constraint that the pseudomodes have physical parameters and include non-Hermitian terms in the pseudomode Hamiltonian or negative damping rates in the GKSL master equation Lambert et al. (2019); Luo et al. (2023); Menczel et al. (2024); Pleasance et al. (2020), or add additional dissipators acting non-locally on the system and pseudomodes Park et al. (2024); Huang et al. (2025), in order to model an even wider variety of environments.
The success of the pseudomodes approach, however, relies on the careful design of their geometry and parameters so as to properly mimic the real environment. Implementing this entails various subtleties, which are often overlooked. The goal of this paper is to address said subtleties. In section II, we consider the effective spectral density that arises from a particular configuration of pseudomodes. Depending on the non-Hermitian Hamiltonian generated by the pseudomodes’ Hamiltonian and the residual baths, three possible cases emerge. If none of the pseudomodes couple to each other, then one obtains an effective spectral density that is a sum of Lorentzians, while if the pseudomodes couple to each other there are additional terms in the effective spectral density which we call “Anti-Lorentzian”. We take this one step further, and show that one can obtain terms which are neither Lorentzian nor Anti-Lorentzian by using a non-diagonalizable non-Hermitian Hamiltonian for the pseudomodes, a case which to the best of our knowledge has not been previously explored.
In section III we discuss the reverse process of choosing pseudomode parameters to obtain a particular effective spectral density, which is required to actually implement the pseudomode method. This is generally done in two steps. First, the true spectral density is fit in time space with a set of decaying complex exponentials. Then, the pseudomode parameters must be chosen to match that exponential fit, which is typically done by imposing some additional structure on the pseudomode parameters and then performing optimization Mascherpa et al. (2020). This second step is the greatest challenge to implementing the pseudomode method, in large part because there are many more free parameters for the pseudomodes than there are constraints given by the effective spectral density. We present a new method for choosing the pseudomode parameters, by which the parameters can be chosen to exactly match a decaying exponential fit. Our method also sheds light on the vast number of free parameters involved when choosing the pseudomode parameters.
@@
In section IV, we briefly explore the use of uncoupled pseudomodes uniformly distributed in energy and show that, contrary to conventional assumption, the effective spectral density obtained from this approach approximates but does not converge to the true bath spectral density. Instead, we present a new, alternate method which , while also not convergent in the infinite limit, significantly reduces the error between the true and effective spectral densities in the limit of many pseudomodes and is equally easy to implement. Finally, in section V we show how the effective spectral densities of section II also emerge in the context of scattering theory for non-interacting systems Blanter and Büttiker (2000); Büttiker (1992); Levitov et al. (1996); Esposito et al. (2009). We confirm that if the pseudomode parameters are chosen such that the effective spectral density exactly matches the true spectral density then the currents obtained from Landauer-Büttiker theory will also match exactly between the original system-bath description and pseudomode description. This also presents a new way to study scattering theory currents by breaking down a current between two real baths into the sum of currents between residual baths under a pseudomode mapping.
@@ I.1 Model System
$H_{\rm tot}=$ $H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k\in\alpha}\omega_{k}d_{\alpha k}^{\dagger}d_{\alpha k}$ (1)
$+\sum_{\alpha\in\{L,R\}}\sum_{i\in S}\sum_{k\in\alpha}\left(g_{\alpha i,k}c_{i}^{\dagger}d_{\alpha k}+g_{\alpha i,k}^{*}d_{\alpha k}^{\dagger}c_{i}\right)\,$,
where $d_{\alpha k}$ is the $k$th fermionic mode of bath $\alpha\in\{L,R\}$. We set $\hbar=1$ here and throughout the paper. The system Hamiltonian can be completely general and may include arbitrary driving. The last term describes how bath $\alpha\in\{L,R\}$ can, in principle, couple to any mode $c_{i}$ of the system. Each bath is characterized by a spectral density matrix
$(J_{\alpha})_{ij}=2\pi\sum_{k\in\alpha}g_{\alpha i,k}^{*}g_{\alpha j,k}\delta(\omega-\omega_{k})\,$, (2)
@@
$\frac{dc_{i}(t)}{dt}=i[H_{S}(t),c_{i}(t)]-i\sum_{\alpha}\sum_{k\in\alpha}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ (3)
$-\sum_{\alpha}\sum_{k\in\alpha}\sum_{j\in S}\intop_{0}^{t}dt^{\prime}g_{\alpha i,k}^{*}g_{\alpha j,k}e^{-i\omega_{\alpha k}(t-t^{\prime})}c_{j}(t^{\prime})\,$,
@@
$\frac{dc_{i}(t)}{dt}$ $=i[H_{S}(t),c_{i}(t)]+\sum_{\alpha}\xi_{\alpha i}(t)$ (4)
$-\sum_{\alpha}\sum_{j\in S}\intop_{0}^{t}dt^{\prime}\chi_{\alpha,ij}(t-t^{\prime})c_{j}(t^{\prime})$ (5)
where $\xi_{\alpha i}=-i\sum_{k\in\alpha}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ is the noise operator, and the memory kernel is given by the Fourier transform of the spectral density:
@@
$C^{+}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)f_{\alpha}(\omega).$ (7)
@@
$\sum_{\alpha}\sum_{i\in S}c_{i}^{\dagger}B_{\alpha i}+B_{\alpha i}^{\dagger}c\,$, (9)
@@
$C^{+}_{\alpha,ij}(t)=\left\langle B_{i}^{\dagger}(t)B_{j}(0)\right\rangle\,$, (10)
$C^{-}_{\alpha,ij}(t)=\left\langle B_{i}(t)B_{j}^{\dagger}(0)\right\rangle\,.$ (11)
@@ I.2 Pseudomode Mapping
$H_{\rm tot}=H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k,q=1}^{n_{\alpha}}\Lambda_{\alpha,kq}a_{\alpha k}^{\dagger}a_{\alpha q}$
$\qquad+\sum_{\alpha\in\{L,R\}}\sum_{i\in S}\sum_{k=1}^{n_{\alpha}}\big(\zeta_{\alpha i,k}a_{\alpha k}^{\dagger}c_{i}+\zeta_{\alpha i,k}^{*}c_{i}^{\dagger}a_{\alpha k}\big)$ (12)
$+\sum_{\alpha\in\{L,R\}}\sum_{k=1}^{n_{\alpha}}\sum_{p\in(\alpha,k)}\Big\{\omega_{p}b_{\alpha k,p}^{\dagger}b_{\alpha k,p}+\tau_{\alpha k,p}\big(a_{\alpha k}^{\dagger}b_{\alpha k,p}+b_{\alpha k,p}^{{\dagger}}a_{\alpha k}\big)\Big\}\,.$
The first line of (I.2) describes the possible interactions between pseudomodes of the same bath, described by the $n_{\alpha}\times n_{\alpha}$ matrix $\Lambda_{\alpha}$ for each bath. If $\Lambda$ is diagonal then none of the pseudomodes couple to each other, and we have the simpler configuration shown in Fig. 1(b). The second line describes how each pseudomode interacts with the central system through a set of $n_{\alpha}\times n_{S}$ matrices $\zeta_{\alpha}$. Finally, the last line describes a set of infinite dimensional residual reservoirs, one for each pseudomode $a_{\alpha k}$, labeled by the pair of indices $(\alpha,k)$ and composed of a set of modes $b_{\alpha k,p}$. The residual baths then have spectral densities given by
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p\in(\alpha,k)}\left|\tau_{\alpha k,p}\right|^{2}\delta(\omega-\omega_{p})\,$, (13)
which is a scalar, rather than a matrix like (2), since each residual bath $(\alpha,k)$ couples only to the pseudomode $a_{\alpha k}$. Generally, we want the couplings to the residual baths $\tau_{\alpha k,p}$ to be constant at all energies, so that the residual baths have constant spectral densities
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha,kk}\,.$ (14)
This allows us to treat the “extended system” consisting of the original system and the pseudomodes using a Markovian master equation. In this case, the (original) system feels an effective memory kernel given by Mascherpa et al. (2020):
@@
The derivation for these forms of $J^{\rm eff}_{\alpha,ij},\chi_{\alpha,ij}^{\rm eff}$ can be found in Appendix A. The pseudomode parameters are then chosen so that the effective spectral density (17) and effective memory kernel (15) are good approximations for the true spectral density (2) and memory kernel (6) of the original system. The reduced system dynamics under (1) will then be identical to those under (I.2) so long as the correlation functions from $J^{\rm eff}_{\alpha,ij}$ are identical to the true correlation functions (7); likewise, the dynamics under (1) will be well-approximated by the dynamics under (I.2) so long as the effective correlation functions well-approximate the true ones Trivedi et al. (2021); Huang et al. (2024). In practice, we achieve this by ensuring that $J^{\rm eff}_{\alpha,ij}(\omega)$ is a good fit to $J_{\alpha,ij}(\omega)$, while setting the temperatures and chemical potentials of the residual baths $(\alpha,k)$ to be the same as those of the original baths $\alpha$.
The utility of this method thus hinges on our ability to choose a set of $\zeta_{\alpha i}$, $\Lambda_{\alpha}$, and $\Gamma_{\alpha}$ so that the effective spectral density (17) closely matches the actual spectral density (2). In the next section, we describe how to obtain the form of $J^{\rm eff}_{\alpha,ij}$ from the chosen pseudomode configuration; in the following section, we describe the more complicated problem of choosing pseudomode parameters to obtain a desired $J^{\rm eff}_{\alpha,ij}$. As we shall show shortly, there are many possible choices of pseudomode parameters that give rise to the same effective spectral density, and not all of them may be equally desirable. Actually computing the extended-system dynamics generally requires further approximations, most commonly the use of a local GKSL master equation, which may be more or less reasonable to make based on the choice of pseudomode parameters. We do not discuss the merits of particular choices of parameters in this paper, instead focusing on the procedure for constructing the pseudomode parameters and emphasizing the freedoms therein. Note also that while in this paper we focus on fermionic systems, the expressions (15) and (17) are identical for bosonic systems under the rotating wave approximation, where the couplings between the system and residual bath take the form (1Mascherpa et al. (2020); Medina et al. (2021), as are the results of Section II, which follow from them. For bosonic systems with position-position coupling (i.e, $(c+c^{\dagger})(d+d^{\dagger})$), one obtains slightly different results, but the general ideas of this paper may remain applicable. Some additional care is required when treating bosonic systems, however, since (17) will typically produce an effective spectral density which is nonzero for $\omega<0$.